How do you find an equation in slop-intercept form of the line satisfying the specified conditions? through (-6,-9), perpendicular to -7x-5y=-3
Answer by stanbon(75887)

How do you find an equation in slop-intercept form of the line satisfying the specified conditions? through (-6,-9), perpendicular to -7x-5y=-3
----------
Find the slope of the given line:
y = (-7/5)x+(3/5)
That slope is -7/5
------------------------------
A perpendicular line must have slope = 5/7
------------------------------------------------
Form of the equation you want:
y = mx +b
You know m = 5/7 and you know y = -9 when x = -6:
-9 = (5/7)(-6) + b
-63/7 = -30/7 + b
b = -33/7
==================
Equation:
y = (5/7)x - (33/7)
===========================
Cheers,
Stan H.